# A RankNet-Inspired Surrogate-Assisted Hybrid Metaheuristic for Expensive Coverage Optimization

### v1

@@ I Introduction
Coverage optimization deals with mathematical models and algorithms together to deploy a set of facilities that satisfy specified customers or demands points in the best possible way. It has an enduring and strong interaction with many other research disciplines including location science [71], sensor networks [27], geography [72], and wireless communication [73].
In the past few decades, coverage optimization not only proved its success in wide applications, but also gained numerous model variants [74]. They are mainly divided into three categories including area coverage, target coverage, and barrier coverage [35]. Driven by modern applications and leveraging structure mathematical optimization, coverage optimization is often formulated as Mixed-Variable Optimization Problems (MVOPs), also known as mixed-integer programming. The discrete variables can represent site selection, which is often described by selecting a subset of candidate sites for deploying facilities (e.g., sensors) [28, 9]. The continuous variables provide more flexibility in modeling emerging issues in coverage optimization. Take an example, directional sensors (e.g., surveillance cameras, radars) have attracting increasing attention in practical applications and introducing additional complexities such as angular parameter optimization [5, 32] and power capacity [29]. MVOPs are rather challenging due to the multiple disconnected regions in its solution space [7]. Moreover, the correlation between discrete and continuous variables is a crucial factor affecting overall performance [5].
@@
Figure 1: A high-fidelity realistic coverage optimization scenario in our previous work [5]. The upper heatmap of Fig. 1a represents the demand distribution, while Fig. 1b and 1c show the distribution of demand points within the deployment area under low and high discretization resolutions, respectively. It is obvious that the computational costs are expensive as the pairs of supply-demand points mount up [33] For example, solving a case with 100 candidate sites, 300 dimensions, and 30,000 targets require over 280 hours.
In this paper, we further rethink coverage optimization via the lens of Expensive Optimization Problems (EOPs) from the following three aspects,
@@
    Target coverage dominates the majority of all literature [74], and is commonly used as an approximation of comprehensive area coverage, with sampling methods employed to discretize the area [75, 32, 63], as shown in Fig. 1. Though finer discretization is more realistic, it significantly increases the complexity of the problem [33].
@@
    Realistic coverage optimization in 2-D or 3-D space calls for Line-of-Sight (LoS) using visibility analysis. Visibility analysis requires incorporating real Digital Elevation Map (DEM) data and performing calculations via interpolation. As the resolution in 1) increases, the pairs of supply and demand points mounts up, and the Euclidean distance between points grows. Consequently, the computational time required for the problem rises significantly [12].
@@
    Evolutionary Algorithms (EA), also known as metaheuristic, have been widely applied to solve coverage optimization [52, 61] and MVOPs due to their robust search capabilities and flexible frameworks [76]. However, inherited difficulties in 1) and 2) pose significant challenges for fitness evaluations (FEs) of individuals in canonical EAs during evolution.
@@
Surrogate-assisted evolutionary algorithms (SAEAs) introduce surrogate models into the EA framework, and have been able to efficiently solve expensive optimization problems (EOPs), such as Computational Fluid Dynamics (CFD) design [64] and Network Architecture Search (NAS)[6]. Therefore, the design of efficient SAEAs has been a key focus in the community. Numerous surrogate models, including Kriging [26], Radial Basis Function Network (RBFN)[7], and Polynomial Response Surface (PRS)[65], have been effectively combined with classical EAs [67]. These models have demonstrated outstanding performance in solving both single and multi-objective optimization problems [68].
Despite the advancements in SAEAs, research on MVOPs remains less studied. Liu et al.[7] and Xie et al. [50] introduce global-local SAEA frameworks, which employ RBFN as surrogate models to estimate the best-performing individual for real fitness evaluation. Similarly, Li et al. [6] combine EDA with Kriging to efficiently solve NAS. Diverging from the evolutionary computation framework, Saves et al. [45] adopt Kriging as the surrogate model to develop a Bayesian-based algorithm tailored for MVOPs.
However, a common limitation of these methods is their reliance on regression-based surrogate models, which struggle to accurately approximate fitness values. This restricts their applicability to MVOPs with fewer than 100 dimensions. As we know, EAs depend on iterative comparisons between parents and offspring for selection [39]. Due to the discontinuous solution space of MVOP, the performance of regression-based surrogates deteriorates with increasing dimension, as their construction depends on distance-based similarity measures [46]. Fortunately, pairwise-based approaches can mitigate this issue by converting continuous value predictions into rankings [40, 42], which enable the models to infer superior individuals for the next generation without accurate fitness function fitting.
Inspired by recommendation systems [17], we propose the RankNet-inspired Surrogate Hybrid Optimization tailored for realistic coverage optimization. As the extension of our previous work [5], RI-SHM leverages the well-established global-local SAEA framework, which has proven to be effective in numerous studies [69, 7, 50]. The RI-SHM contains three key components: (1) a RankNet-based pairwise global surrogate model, (2) a surrogate-assisted EDA, and (3) an efficient switching strategy.
@@
    We reformulate coverage optimization via the lens of EOPs to create a high-fidelity realistic model, and propose an efficient RI-SHM to address this EMVOP.
@@
    We construct a RankNet-based pairwise global surrogate to predict the rankings between pairs of individuals. To the best of our knowledge, this is the first application of a pairwise-based surrogate model to MVOPs, which is realized within the context of modern deep learning techniques. In addition, we integrate RBFN with EDA to refine local regions and propose a fitness diversity-driven strategy for adaptively switching between global and local optimization phases.
@@ II Related work
In this section, we first review the related coverage optimization models for target coverage. Then, we explore recent advancements in surrogate models within SAEAs, focusing primarily on single-objective optimization problems while referencing studies on other problems as well.
@@ II-A Coverage Optimization
With the rise of sensor networks, coverage optimization has gained increasing attention which leads to numerous model variants [74]. They are generally divided into three categories including area coverage, target coverage, and barrier coverage [35]. Among them, target coverage dominates the majority of the literature, where the area is discretized through sampling methods to approximate comprehensive area coverage [32, 75, 63]. In this part, we focus on target coverage for sensor networks.
@@
Other researchers have approached coverage optimization as MVOPs. These models extend beyond site selection, integrating emerging issues such as power capacity [29, 10] to address real-world constraints. Zhu et al. [29, 10, 41] investigate the deployment of visual sensor networks in 3-D space, aiming to maximize target coverage while minimizing power consumption. Specifically, the sensor’s working direction and charging period are optimized as continuous variables. Another study by Nguyen et al. [41] incorporates cluster heads alongside sensor nodes to balance energy consumption for precision farming, which treats energy consumption as continuous variables.
With the increasing adoption of directional sensors in practical applications, we model our coverage optimization as a MVOP that simultaneously optimizes both site selection and angular parameters [9, 32, 5]. To ensure variable consistency, previous studies often address the mixed variables by discretizing continuous variables. Saad et al. [9] aim to maximize target coverage while minimizing the number of sensors. They discretize continuous angles and integrate them with site selection to formulate a binary-encoded Multi-Objective Problem (MOP). Similarly, Heyn [32] adopts a coarser approach by discretizing angles into several spatial poses to formulate a discrete optimization problem for backup coverage. However, the simplification of discretizing continuous angles may lead to the loss of precision.
To preserve the precision, we proposed $CAGA_{mv}$ that effectively captures the correlation between mixed variables in our previous work [5]. As mentioned before, using target coverage to approximate comprehensive area coverage is commonplace. However, the optimal deployment solution is highly sensitive to the level of discretization. While finer discretization provides a more realistic model, it significantly increases the complexity of our problem [33]. Moreover, realistic coverage optimization requires visibility analysis to obtain LoS. As the pairs of supply and demand points mount up, the computational cost grows substantially [12]. This issue becomes even more severe for population-based EAs [53]. To address these issues, we introduce SAEAs for more efficient solutions.
@@ II-B Surrogate Models in SAEAs
Regression-based models are the most widely used surrogates in SAEAs. They facilitate environment selection by directly predicting the fitness values of individuals. Liu et al. [46] introduce GPEME, a classic SAEA that utilizes Kriging for medium-scale problems, while Zhou et al. [47] apply Kriging to optimize parameter identification of battery models in real world. However, Kriging’s training complexity is $O(N^{3}\cdot D)$ [46], where $N$ is the number of training samples and $D$ stands for the dimension, which becomes computationally prohibitive for large-scale problems. In contrast, RBFNs are favored for their simplicity and lower computational cost [50]. Liu et al. [7] propose SHEALED, which employs a global-local framework with RBFNs as surrogates for high-dimensional MVOPs.
Classification-based models are often used to predict the qualitative results between individuals and reference solutions, aligning closely with the selection process in EAs. Lu et al. [44] employ RankSVM as a classifier to identify the most promising trial vector in Differential Evolution (DE). For MOPs, Zhang et al. [49] train a $k$-Nearest Neighbor (kNN) classifier to distinguish non-dominated solutions from dominated ones. Similarly, CSEA [48] utilize a Deep Neural Network (DNN) to determine whether individuals dominate predefined reference points.
Recently, pairwise-based models have gained attention. Unlike the classification-based ones, pairwise-based surrogates focus on determining the relative preference between arbitrary pairs of individuals, eliminating the need for absolute definitions of good or bad solutions [38]. Dushatskiy et al. [43] propose the first Convolutional Neural Network (CNN) -based surrogate capable of predicting the fitness relationship between two individuals, and validate it on Binary Optimization Problems (BOPs). However, most pairwise-based surrogates are tailored for predicting Pareto dominance [40, 39, 42], leaving a notable gap in their application to MVOPs.
Table. I briefly concludes the common surrogate models used in SAEAs. MVOPs pose significant challenges due to the discontinuity and complexity of the solution space [7], making it difficult for regression-based models to accurately predict fitness values. Pairwise-based models offer greater robustness than regression-based ones by focusing on relative ordering rather than precise value predictions. Furthermore, advancements in deep learning provide us great opportunity to effectively model the mixed-variable correlations, paving the way for efficient solutions to large-scale MVOPs.
@@ III-A Collaborative Coverage Model
Our coverage optimization involves selecting a subset of $k$ sites from $|Z|$ candidates to deploy directional sensors while simultaneously determining their corresponding angle parameters. Consequently, each target $q$ is collaboratively covered by a network of multiple sensors, with the coverage probability of $q$ defined as:
@@
The sensor network $S$ is required to cover a set of target points $Q$. To facilitate customized surveillance, each target $q\in Q$ is assigned with a weight $\omega_{q}$. Targets closer to critical areas will receive higher weights. Our objective is to maximize coverage of critical targets while ensuring broad coverage of others as well. To summarize, we formulate the mathematical model as follows
@@ III-B Probabilistic Sensing Model
Recent research has shifted from traditional binary and omnidirectional models to directional and probabilistic ones, which better address real-world uncertainties. In 3-D space, coverage is further constrained by terrain obstructions, which requires visibility analysis that often involves costly interpolation. Formally, the coverage probability is defined as,
@@
Coverage of target $q$ by sensor $s_{i}$ requires $q$ to lie within $s_{i}$’s detection range; in addition, the visibility (i.e., Line-of-Sight, LoS) between them should be ensured. We employ the Bresenham LoS algorithm [11] to analyze the visibility between two points and represent the presence of visibility using a binary variable $v(s_{i},q)\in\{0,1\}$.
@@ III-C Discussion
Furthermore, Eq. 3 indicates that realistic coverage optimization requires visibility analysis $v(s_{i},q)$, which relies on real DEM data and interpolation-based calculations. As the resolution of discretization increases, both the number of supply-demand point pairs $(s_{i},q)$ and the Euclidean distance between points grow. As a result, the corresponding computational cost arises significantly [12]. This explains why high-fidelity realistic coverage optimization evolve into EMVOP challenges.
The issues outlined above poses a significant challenge for canonical EAs, as the evaluation of each individual’s fitness greatly increases the computational burden [13]. To address this, we propose a novel SAEA inspired by RankNet to enhance evolutionary efficiency in the next section.
Figure 2: The generic framework of the proposed RI-SHM, which consists of three key components: a RankNet-based pairwise global surrogate model, a surrogate-assisted local EDA, and a fitness diversity-driven switching strategy.
@@ IV-A Overall Framework
The high-fidelity realistic coverage optimization of sensor networks, classified as an EMVOP, can benefit from surrogate-assisted optimization to enhance both efficiency and performance. To this end, we introduce the RI-SHM within a global-local SAEA framework, which consists of three key components: a RankNet-based pairwise global surrogate, a surrogate-assisted local EDA, and a fitness diversity-driven switching strategy. The generic framework of our RI-SHM is shown in Fig. 2, which mainly includes the following five steps:

### v2

@@ I Introduction
Coverage optimization deals with mathematical models and algorithms to deploy a set of facilities that satisfy specified customers or demand points in the best possible way. It has an enduring and intense interaction with many other research disciplines, including location science [1], sensor networks [2], geography [3], and wireless communication [4].
In the past few decades, coverage optimization proved its success in wide applications and gained numerous model variants [5]. Driven by modern applications and leveraging structure mathematical optimization, coverage optimization is often formulated as Mixed-Variable Optimization Problems (MVOPs), also known as mixed-integer programming. The discrete variables can represent site selection, which is often described by selecting a subset of candidate sites for deploying facilities (e.g., sensors) [6, 7]. The continuous variables provide more flexibility in modeling emerging issues in coverage optimization. Take an example: directional sensors (e.g., surveillance cameras, radars) have attracted increasing attention in practical applications and introducing additional complexities such as angular parameter optimization [8, 9] and power capacity [10]. MVOPs are challenging due to the multiple disconnected regions in its solution space [11]. Moreover, the correlation between discrete and continuous variables is a crucial factor affecting overall performance [8].
@@
Figure 1: A high-fidelity realistic coverage optimization scenario in our previous work [8]. The upper heatmap of Fig. 1a represents the demand distribution, while Fig. 1b and 1c show the distribution of demand points within the deployment area under low and high discretization resolutions, respectively. It is obvious that the computational costs are expensive as the pairs of supply-demand points mount up [12]. For example, solving a case with 100 candidate sites, 300 dimensions, and 30,000 targets requires over 280 hours.
Moreover, target coverage in coverage optimization can be effectively modeled via the lens of Expensive Optimization Problems (EOPs) due to the following three aspects,
@@
    Target coverage dominates the majority of all literature [5], and is commonly used as an approximation of comprehensive area coverage, with sampling methods employed to discretize the area [13, 9, 14], as shown in Fig. 1. Though finer discretization is more realistic, it significantly increases the problem complexity [12].
@@
    Realistic coverage optimization in 2-D or 3-D space calls for Line-of-Sight (LoS) calculation with visibility analysis. Visibility analysis requires incorporating real Digital Elevation Map (DEM) data and performing calculations via interpolation. As the resolution in 1) increases, the supply and demand point pairs mount, and the Euclidean distance between points grows. Consequently, the computational time required for the problem rises significantly [15].
@@
    Evolutionary Algorithms (EA), also known as metaheuristics, have been widely applied to solve coverage optimization [16] and MVOPs due to their robust search capabilities and flexible frameworks [17]. However, inherited difficulties in 1) and 2) pose significant challenges for individuals’ fitness evaluations (FEs) in canonical EAs during evolution.
@@
Surrogate-assisted evolutionary algorithms (SAEAs) introduce surrogate models into the EA framework and have been able to efficiently solve EOPs, such as Computational Fluid Dynamics design and Network Architecture Search (NAS)[18]. Therefore, the design of efficient SAEAs has been a key focus in the community. Numerous surrogate models, including Kriging [19], Radial Basis Function Network (RBFN)[11], and Polynomial Response Surface, have been effectively combined with classical EAs [20]. These models have demonstrated outstanding performance in solving single and multi-objective optimization problems [21].
Despite the advancements in SAEAs, research on MVOPs remains less studied. Liu et al.[11] and Xie et al. [22] introduce global-local SAEA frameworks, which employ RBFN as surrogate models to estimate the best-performing individual for real fitness evaluation. Similarly, Li et al. [18] combine the Estimation of Distribution Algorithm (EDA) with Kriging to efficiently solve NAS. Diverging from the evolutionary computation framework, Saves et al. [23] adopt Kriging as the surrogate model to develop a Bayesian-based algorithm tailored for MVOPs.
However, a common limitation of these methods is their reliance on regression-based surrogate models, which struggle to approximate fitness values accurately. This restricts their applicability to MVOPs with fewer than 100 dimensions. As we know, EAs depend on iterative comparisons between parents and offspring for selection [24]. Due to the discontinuous solution space of MVOP, the performance of regression-based surrogates deteriorates with increasing dimension, as their construction depends on distance-based similarity measures [25]. Fortunately, pairwise-based approaches can mitigate this issue by converting continuous value predictions into rankings [26, 27], which enables the models to infer superior individuals for the next generation without accurate fitness function fitting.
Inspired by recommendation systems [28], we propose the RankNet-inspired Surrogate Hybrid Optimization tailored for realistic coverage optimization. As the extension of our previous work [8], RI-SHM leverages the well-established global-local SAEA framework, which has proven to be effective in numerous studies [11, 22]. The RI-SHM contains three key components: (1) a RankNet-based pairwise global surrogate model, (2) a surrogate-assisted EDA, and (3) an efficient switching strategy.
@@
    We reformulate coverage optimization via the lens of EOPs to create a high-fidelity realistic model and propose an efficient RI-SHM to address this EMVOP.
@@
    We construct a RankNet-based pairwise global surrogate to predict the rankings between pairs of individuals. Notably, this represents a novel adaptation of pairwise-based surrogate models for MVOPs implemented within modern deep-learning techniques. In addition, we integrate RBFN with EDA to refine local regions and propose a fitness diversity-driven strategy for adaptively switching between global and local optimization phases.
@@ II Related work
In this section, we first review the related coverage optimization models for target coverage. Then, we explore recent advancements in surrogate models within SAEAs, focusing primarily on single-objective optimization problems and referencing studies on other problems.
@@ II-A Coverage Optimization
With the rise of sensor networks, coverage optimization has gained increasing attention, which leads to numerous model variants [5]. They can be divided into three main categories: area coverage, target coverage, and barrier coverage [34]. Among them, target coverage dominates the majority of the literature, where the area is discretized through sampling methods to approximate comprehensive area coverage [9, 13, 14]. In this part, we focus on target coverage for sensor networks.
@@
Other researchers have approached coverage optimization as MVOPs. These models extend beyond site selection, integrating emerging issues such as power capacity [36] to address real-world constraints. Zhu et al. [36] investigate the deployment of visual sensor networks in 3-D space, aiming to maximize target coverage while minimizing power consumption. Specifically, the sensor’s working direction and charging period are optimized as continuous variables. Another study by Nguyen et al. [37] incorporates cluster heads alongside sensor nodes to balance energy consumption for precision farming, which treats energy consumption as continuous variables.
With the increasing adoption of directional sensors in practical applications, we model our coverage optimization as an MVOP that simultaneously optimizes both site selection and angular parameters [7, 9, 8]. To ensure variable consistency, previous studies often address the mixed variables by discretizing continuous variables. Saad et al. [7] aims to maximize target coverage while minimizing the number of sensors. They discretize continuous angles and integrate them with site selection to formulate a binary-encoded Multi-Objective Problem (MOP). Similarly, Heyn [9] adopts a coarser approach by discretizing angles into several spatial poses to formulate a discrete optimization problem for backup coverage. However, simplifying discretizing continuous angles may lead to the loss of precision.
To preserve the precision, we proposed $CAGA_{mv}$ that effectively captures the correlation between mixed variables in our previous work [8]. As mentioned before, using target coverage to approximate comprehensive area coverage is commonplace. However, the optimal deployment solution is highly sensitive to the discretization level. While finer discretization provides a more realistic model, it significantly increases the complexity of our problem [12]. Moreover, realistic coverage optimization requires visibility analysis to obtain LoS. As the pairs of supply and demand points mount up, the computational cost grows substantially [15]. This issue becomes even more severe for population-based EAs [38]. To address these issues, we introduce SAEAs for more efficient solutions.
@@ II-B Surrogate Models in SAEAs
Regression-based models are the most widely used surrogates in SAEAs. They facilitate environment selection by directly predicting the fitness values of individuals. Liu et al. [25] introduce GPEME, a classic SAEA that utilizes Kriging for medium-scale problems, while Zhang et al. [29] apply Kriging to optimize energy absorption of multi-cell structures in the real world. Moreover, Saves et al. [23] extend Kriging to a mixed-variable version, enabling its application to MVOPs. However, Kriging’s training complexity is $O(N^{3}\cdot D)$ [25], where $N$ is the number of training samples and $D$ stands for the dimension, which becomes computationally prohibitive for large-scale problems. In contrast, RBFNs are favored for their simplicity and lower computational cost [22]. Liu et al. [11] propose SHEALED, which employs a global-local framework with RBFNs as surrogates for high-dimensional MVOPs.
Classification-based models often predict the qualitative results between individuals and reference solutions, aligning closely with the selection process in EAs [41]. Lu et al. [30] employ RankSVM as a classifier to identify the most promising trial vector in Differential Evolution (DE). For MOPs, Zhang et al. [31] train a $k$-Nearest Neighbor (kNN) classifier to distinguish non-dominated solutions from dominated ones. Similarly, CSEA [32] utilizes a Deep Neural Network (DNN) to determine whether individuals dominate predefined reference points.
Recently, pairwise-based models have gained attention. Unlike classification-based surrogates, pairwise-based surrogates focus on determining the relative preference between arbitrary pairs of individuals, eliminating the need for absolute definitions of good or bad solutions [26]. Dushatskiy et al. [33] propose the first Convolutional Neural Network (CNN) -based surrogate capable of predicting the fitness relationship between two individuals and validating it on Binary Optimization Problems (BOPs). However, most pairwise-based surrogates are tailored for predicting Pareto dominance [26, 24, 27], leaving a notable gap in their application to MVOPs.
Table. I briefly concludes the common surrogate models used in SAEAs, highlighting the limited research on efficient algorithms for expensive MVOPs. MVOPs pose significant challenges due to the discontinuity and complexity of the solution space [11], making it difficult for regression-based models to predict fitness values accurately. Pairwise-based models offer greater robustness than regression-based ones by focusing on relative ordering rather than precise value predictions. Furthermore, advancements in deep learning provide us with a great opportunity to effectively model mixed-variable correlations, paving the way for efficient solutions to large-scale MVOPs.
@@ III-A Collaborative Coverage Model
Our coverage optimization involves selecting a subset of $k$ sites from $|Z|$ candidates to deploy directional sensors while determining their corresponding angle parameters. Consequently, each target $q$ is collaboratively covered by a network of multiple sensors, with the coverage probability of $q$ defined as:
@@
The sensor network $S$ is required to cover a set of target points $Q$. To facilitate customized surveillance, each target $q\in Q$ is assigned with a weight $\omega_{q}$. Targets closer to critical areas will receive higher weights. We aim to maximize coverage of critical targets while ensuring broad coverage of others. To summarize, we formulate the mathematical model as follows
@@ III-B Probabilistic Sensing Model
Recent research has shifted from traditional binary and omnidirectional models to directional and probabilistic ones, which better address real-world uncertainties. In 3-D space, terrain obstructions further constrain coverage, which requires visibility analysis that often involves costly interpolation. Formally, the coverage probability is defined as,
@@
Coverage of target $q$ by sensor $s_{i}$ requires $q$ to lie within $s_{i}$’s detection range; in addition, the visibility (i.e., Line-of-Sight, LoS) between them should be ensured. We employ the Bresenham LoS algorithm [42] to analyze the visibility between two points and represent the presence of visibility using a binary variable $v(s_{i},q)\in\{0,1\}$. Although the Bresenham LoS algorithm avoids extensive interpolation calculations, it becomes computationally expensive in high-dimensional, large-scale scenarios. For instance, solving a problem with 100 candidate sites, 300 dimensions, and 30,000 targets requires over 280 hours.
@@ III-C Discussion
Furthermore, Eq. 3 indicates that realistic coverage optimization requires visibility analysis $v(s_{i},q)$, which relies on real DEM data. As the resolution of discretization increases, both the number of supply-demand point pairs $(s_{i},q)$ and the Euclidean distance between points grow. As a result, the corresponding computational cost arises significantly [15]. This explains why high-fidelity realistic coverage optimization evolves into EMVOP challenges.
The issues outlined above pose a significant challenge for canonical EAs, as the evaluation of each individual’s fitness dramatically increases the computational burden [39]. To address this, we propose a novel SAEA inspired by RankNet to enhance evolutionary efficiency in the next section.
Figure 2: The generic framework of the proposed RI-SHM consists of three key components: a RankNet-based pairwise global surrogate model, a surrogate-assisted local EDA, and a fitness diversity-driven switching strategy.
@@ IV-A Overall Framework
The high-fidelity realistic coverage optimization of sensor networks, classified as an EMVOP, can benefit from surrogate-assisted optimization to enhance efficiency and performance. To this end, we introduce the RI-SHM within a global-local SAEA framework, which consists of three key components: a RankNet-based pairwise global surrogate, a surrogate-assisted local EDA, and a fitness diversity-driven switching strategy. The generic framework of our RI-SHM is shown in Fig. 2, which mainly includes the following five steps: